If one-half of an interger is subtracted from three-fifths of the next consecutive interger, the difference is 3. What are the two intergers?
I am lost on trying to figure out this problem, I can solve them without the fractions but with confuse me horribly. Thank you for the help

Answer by checkley71(8403)

(3/5)(X+1)-X/2)=3
(3X+3)/5-X/2=3
[(3X+3)*2-5X]/10=3
(6X+6-5X)/10=3
X+6=30
X=30-6
X=24
PROOF
(3/5)(24+1)-24/2=3
3/5*25-12=3
75/5-12=3
15-12=3
3=3
###########################################
MERRY CHRISTMAS